Please show me how to solve the equation using the Zero-factor property. Thanks!
X^2+4x-21=0

Answer by tutorcecilia(2152)

X^2+4x-21=0
(x+7)(x-3)=0 [factor]
x+7=0 [set each factor equal to zero and solve]
x=-7
and
x-3=0
x=3
.
check by plugging in each value for x and solve.